OpenStudy (anonymous):

what is 1 and 4/5 + 2 and 1/3

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

First you need to find a common denominator between 4/5 and 1 / 3 What is the common denominator between those two fractions?

OpenStudy (anonymous):

To turn mixed fractions into regular fractions use this technique: \[1\frac{ 4 }{ 5 }=\frac{ 5*1+4 }{ 5 }=\frac{ 9 }{ 5 }\]

OpenStudy (anonymous):

Once you have done this, find a common denominator and add the two new fractions you have obtained together.

OpenStudy (anonymous):

the lesson said that the answer was 4 and 2/15 I didn't get where they came up with that answer

OpenStudy (anonymous):

http://www.calculatorsoup.com/calculators/math/mixednumbers.php go here and put in the numbers

OpenStudy (anonymous):

Thank you cody77445. Calculator soup was a huge help they showed the way i learned in grade school, my daughters lesson is wacko when it comes to the specifics calculator soup is two thumbs up definitely.

OpenStudy (anonymous):

your welcome it always helps me when i need help